Derive the OLS estimator \hat{β}₀ in the regression model yi=β₀+ui. Show all of the steps in your derivation.
Answer #1

To drive this estimator, we will use the method of moments as it does not use a shred of calculus. In this method, we can identify one parameter for every constraint that we impose on the data.

Firstly we assume our model, then use that assumption as a rule. We make it accurate in our data instead of just taking it as true. The data that we have to make them obey the law. Each rule we impose gives us the ability to say what one parameter must be for the government to hold.

Univariate Regression Model or a model with no \(\mathrm{x}\)-variable: A model where \(\mathrm{y}\) is always equal to constant

$$ y=\beta 0+\mu $$

\(\beta 0\) is umknown parameter in thie model. Therefore we need one assumption \(-\) one rule \(-\) to identify \(\beta 0\). We will impose this rule on the data, and this will give us the formula for our best estimate of \(\beta 0\), which we will denote \(\widehat{\beta 0}\).

The standard assumption we will use is \(E(\mu)=0\), the assumption that the model is correct, on average.

This method works by imposing this assumption on the data. So we start with

$$ E(\mu)=0 $$

and we make this true in the sample of data that we have.

$$ \mu=y-\beta 0 $$

So that

$$ E(\mu)=E(y-\beta 0) $$

It means taking the concept of "expectation," which is a concept about the universe of data, and applying the analogous concept to our sample of data. The mean or average is the sample analog of the expectation. The sampled analog of the expected value of something is the mean of that same thing.

So if the theoretical expected value of a variable \(x\) is \(\mu x\), then the sample analog of \(\mu x\) is \(\bar{x}\). Making \(E(\mu)=0\) true in our data translates to forcing the average of \(u\) in our data to equal zero.

Putting this notion in equation (1), we get

$$ \begin{aligned} &E(\mu)=0=E(y-\beta 0) \\ &=E(y)-E(\beta 0) \\ &=1 / N \sum y i-1 / N \sum \beta 0(2) \\ &=\bar{y}-1 / N * N \beta 0 \\ &=\bar{y}-\beta 0 \text { (3) } \end{aligned} $$

We now have a single, simple equation with a single unknown. We choose our estimate of \(\beta 0\), which we call \(\widehat{\beta 0}\), to be the estimate that solves equation (3).

$$ \begin{aligned} &0=\bar{y}-\widehat{\beta 0}(4) \\ &\widehat{\beta 0}=\bar{y}(5) \end{aligned} $$

The best estimate of a variable \(y\) that we can manage when we model \(y\) as a constant is \(\widehat{\beta 0}=\bar{y}\), the mean of \(y\).